I need help on this question. I know the answer but i dont understand how to get to it...thanks if you could reply this
THE QUESTION IS:
When i was 14 my mother was 41, and she is now twice as old as am. how old am i?
the answer is 27 i know but how do u get to it...thanks, RACHEL

Found 2 solutions by amoresroy, Alan3354:
Answer by amoresroy(361)

Let x = age of the child
x+27 = age of the mother (27=41-14)
X+27 = 2X
2X-X = 27
X = 27

Answer by Alan3354(69443)

The age difference will always be 27.
For the mother to be 2x, it has to be 27 & 54.